OpenStudy (anonymous):

Solve 4x - 2x - 2 = -42.

Parth (parthkohli):

Why not try to first combine the terms, then isolate \(x\)?

OpenStudy (theviper):

4x-2x-2=-42 2x-2=-42 2x=-42+2 2x=-40 x=-40/2=??

OpenStudy (anonymous):

cause im not good at math thats why.

Parth (parthkohli):

\[ \color{Black}{\Rightarrow 4x - 2x - 2 = -42 \Longrightarrow 2x - 2 = 42}\] etc

OpenStudy (anonymous):

2x-2=-42 2x=-42+2 = -40 x=-40/2 =-20

OpenStudy (anonymous):

ty veerashniksai

OpenStudy (anonymous):

k